mathslover (mathslover):

Integral Calculus .... Find the value of integral : \( I = \int \left( \cfrac{x^2}{1+x^6} dx \right) \)

mathslover (mathslover):

@ganeshie8 @BSwan @UnkleRhaukus

OpenStudy (kainui):

x^3=tan(theta)

mathslover (mathslover):

I don't get it that why do we do this? Is there any specific rule for this?

Parth (parthkohli):

http://i.imgur.com/11hGmBW.png

Parth (parthkohli):

The green rectangle.

mathslover (mathslover):

Do you have an integral?

OpenStudy (kainui):

My first attempt was just to set x^2=u, but I realized that wouldn't work. Usually when I see something similar to \[1\ + x^2\] I try to see if I can fit it to a trig substitution. In this case, \[\large x^6=(x^3)^2\] so I just tried it out. It also seems to make sense that this is right because the derivative of x^3 is x^2.

mathslover (mathslover):

^ derivative is 3x^2... And okay, so, the next step will to replace dx by d(tan(theta)) ?

OpenStudy (kainui):

Well I didn't mean that literally, I just meant that x^3 has a scalar multiple derivative of x^2 but whatever lol Yeah. Just take it through like a normal substitution now.

mathslover (mathslover):

That comes out to be : I = \(\int \left(\cfrac{1}{3(1+\tan^2 \theta )} d\tan \theta\right) \)

mathslover (mathslover):

\(d(\tan \theta) =3x^2 dx\) \(dx = \cfrac{d(\tan \theta)}{3x^2}\) So, I just plugged this in at the place of dx in the integral and I got : \(\color{blue}{\text{Originally Posted by}}\) @mathslover That comes out to be : I = \(\int \left(\cfrac{1}{3(1+\tan^2 \theta )} d\tan \theta\right) \) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (kainui):

\[\large x^3=\tan \theta \\ \large 3x^2dx = \sec^2 \theta d \theta\] I'd write it more like this I think.

mathslover (mathslover):

\( I = \cfrac{1}{3} \int (\cos^2 \theta )d\tan \theta \)

mathslover (mathslover):

Hmm well yes, youright... \(I = \cfrac{1}{3} (\cos^2 \theta) \sec^2 \theta d\theta\) Right?

OpenStudy (kainui):

You got it. =)

mathslover (mathslover):

Yes, but I missed integral sign there...

OpenStudy (kainui):

Yeah but you left the dTheta in there so I assumed you knew it wasn't over lol.

mathslover (mathslover):

As the answer comes out to be \(\cfrac{\theta}{3}\) So, finally : \(\cfrac{1}{3} \times \left(\tan^{-1} (x^3) \right) + C\) That should be the final answer, right?

OpenStudy (kainui):

Yep. It sorta worked out much nicer than expected haha.

mathslover (mathslover):

Haha yeah... Thanks for the help.

OpenStudy (kainui):

Yeah, glad I could help. =)

OpenStudy (shamim):

let x^3=z

mathslover (mathslover):

Yeah shamim, I just noticed that we can do it like that too : Let \(x^3 = u\) du = \(3x^2 dx\) \(dx = \cfrac{du}{3x^2}\) That becomes : \(I = \int \cfrac{x^2}{1+(x^3)^2} dx = \int \cfrac{x^2}{1 + u^2} \times \cfrac{du}{3x^2} \) or : \(\int \cfrac{du}{(1+u^2)3} \) Since : \(\int \cfrac{1}{1+x^2} \times dx\) = \(\tan^{-1} x + C\) Therefore, I = \(\cfrac{1}{3} \tan^{-1} (x^3) + C \)

mathslover (mathslover):

Though, @Kainui 's method was also nice. :-)

OpenStudy (kainui):

The inverse trig integrals are just a special case of trig substitution anyways. =P